# Enhancing Sequential Recommendations through Multi-Perspective Reflections and Iteration

### v1

@@ 3.3. Multi-Perspective Reflections Generation
As illustrated in the left of Fig. 2, MoRE aims at exploiting users’ historical interaction to generate reflections from the perspectives of explicit preference, implicit preference, and collaborative signals. We achieve this goal through the following reflectors:
Explicit Preference Reflector infers user $u$’s explicit preferences (e.g., item titles) by driving a LLM to reflect the interacted items $\mathcal{S}_{u}$, the candidate items $\mathcal{C}_{u}$, and the difference between LLM’s prediction $\hat{\mathcal{O}}_{u}$ and ground-truth $\mathcal{O}_{u}$. LLM generate these reflections simultaneously and return the overall reflection $Ref_{\mathrm{EP}}$, which can be formulated as:
(3) $Ref_{\mathrm{EP}}=\mathrm{LLM}_{\mathrm{REF}}(P_{\mathrm{EP}}(u,\mathcal{S}_{u},\mathcal{C}_{u},\hat{\mathcal{O}}_{u},\mathcal{O}_{u})),$
where $\mathrm{LLM}_{\mathrm{REF}}$ is a frozen LLM specific for reflection generation, and $P_{\mathrm{EP}}$ denotes the prompt for reflecting from the perspective of explicit user preferences, which represents items using their titles. The details are shown in App. B.3.
Implicit Preference Reflector is designed to capture the preference embedded in the other item attributes (e.g., item brands) from the interaction sequences $\mathcal{S}_{u}$, the candidate set $\mathcal{C}_{u}$ and the LLM prediction $\hat{\mathcal{O}}_{u}$ and label list $\mathcal{O}_{u}$. This reflector functions similarly to the explicit preference reflector and adopts a frozen LLM to generate reflections:
(4) $Ref_{\mathrm{IP}}=\mathrm{LLM}_{\mathrm{REF}}(P_{\mathrm{IP}}(u,\mathcal{S}_{u},\mathcal{C}_{u},\hat{\mathcal{O}}_{u},\mathcal{O}_{u})),$
where $P_{\mathrm{IP}}$ denotes the reflection prompt from the perspective of implicit user preferences, representing items using their brand names, as detailed in App. B.4.
Collaborative Filtering Reflector targets guiding LLM to reflect from the perspectives of collaborative filtering information. To incorporate the CF signals, instead of introducing new token embeddings into the vocabulary (Zheng et al., 2024) or inserting special sequences into prompts (Zhang et al., 2024), we adopt a pre-trained CF model to rate the CF scores $\mathcal{M}_{u}^{\mathcal{S}}$ and $\mathcal{M}_{u}^{\mathcal{C}}$ respectively for the interacted items for $\mathcal{S}_{u}$ and the candidate items $\mathcal{C}_{u}$. These scores are then involved in the input of the reflector, which is formulated as:
(5) $Ref_{\mathrm{CF}}=\mathrm{LLM}_{\mathrm{REF}}(P_{\mathrm{CF}}(u,\mathcal{S}_{u},\mathcal{C}_{u},\mathcal{M}_{u}^{\mathcal{S}},\mathcal{M}_{u}^{\mathcal{C}},\hat{\mathcal{O}}_{u},\mathcal{O}_{u})),$
where $P_{\mathrm{CF}}$ denotes the reflection prompt from the perspective of CF, as detailed in App. B.5.
@@
We maintain three memory banks for each user to store the generated reflections from the corresponding three perspectives. These memory banks are guided by the meta-reflector to improve the reflections, as will be elaborated in the following section.
3.4. Reflection Memory Maintenance
@@
Reflection Improvement Effect is defined as the difference in LLM recommendation performance on the validation set, with or without the reflections. In other words, if incorporating reflections improves the LLM’s performance, it indicates that the reflection is effective for recommendation. The greater the performance improvement, the more effective the reflection. If the reflection improvement effect of a reflection exceeds a preset threshold, we regard it as an effective one. Formally, we formulate the reflection improvement effect as:
(6) $Imp$ $=\mathrm{Metric}(\hat{\mathcal{O}}_{u}^{\mathrm{REF}})-\mathrm{Metric}(\hat{\mathcal{O}}_{u}^{\mathrm{REC}})$
$\hat{\mathcal{O}}_{u}^{\mathrm{REF}}$ $=\mathrm{LLM}_{\mathrm{REC}}(Ref_{i}(u,\mathcal{S}_{u},\mathcal{C}_{u})),$
$\hat{\mathcal{O}}_{u}^{\mathrm{REC}}$ $=\mathrm{LLM}_{\mathrm{REC}}(P(u,\mathcal{S}_{u},\mathcal{C}_{u})),$
where $Ref_{i}\in\{Ref_{\mathrm{EP}},Ref_{\mathrm{IP}},Ref_{\mathrm{CF}}\}$, $\mathrm{Metric}$ denotes the recommendation metric, e.g., NDCG@10.
@@
(2) Group Level: We use a collaborative filtering (CF) model to construct embeddings for all users and then cluster them using the K-means++ algorithm (Bahmani et al., 2012). We then choose the reflections that provide the greatest average improvement for users within the same cluster, ensuring that users in the same cluster share the same reflections.
@@
All three of the aforementioned strategies are greedy, which may lead to local optima. To mitigate this issue, we propose using the improvement effect brought about by each reflection as the sampling probability. Reflections sampled from the filtered set serve as demonstrations and are exploited in the subsequent iteration step.
@@
3.5. Reflection Selection
@@
The second objective of the meta-reflector is to model preference shifts in SeqRec. Since we have designed three memory banks for each user to collect effective reflections from three different perspectives, we require the meta-reflector to select one of these reflections for recommendation to achieve this goal. We formulate the decision-making process as a Multi-Armed Contextual Bandit (MACB) problem (Li et al., 2010), and adopt the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017) to efficiently learn to select the reflections. Formally, we define MACB as a tuple $(\mathcal{Z},\mathcal{A},\mathcal{R})$, where $\mathcal{Z},\mathcal{A},\mathcal{R}$ denote the state space, action (arm) space, and reward function, respectively.
@@
  • State $\bm{z}\in\mathcal{Z}$ is defined as the recommendation context. We represent it by concatenating the embedding of user $u$ and the interaction sequence $\mathcal{S}_{u}$, i.e., $\bm{z}=[emb_{u},emb_{v}]$, where the embeddings are derived from a CF model, and $emb_{v}$ denotes the average embedding for $\mathcal{S}_{u}$.
@@
  • Action $\bm{a}\in\mathcal{A}$ is defined as the selection of a reflection from three memory banks. We represent $\bm{a}$ as a 3-dimensional vector, with each dimension corresponding to an arm, each arm representing a reflection memory.
@@
  • Reward is defined to assess the recommendation improvement brought by the reflections. We reuse the measure in Eq.(6) and represent the reward as $R(\bm{z,a})=Imp$. Since the reflection has been assessed in Sec. 3.3 we do not need to call the LLM here. Additionally, due to the sampling discussed in Sec. 3.4, $R(\bm{z,a})$ can be both positive and negative, ensuring training efficiency.
@@
  • Replay Buffer is designed to facilitate efficiency in policy optimization. We represent it as $D=(\bm{z},\bm{a},R(\bm{z},\bm{a}),\bm{z}^{\prime})$ to store the tuples of observed state, action, reward, and next state. With the records in the replay buffer, we can refine successful policies and learn from erroneous trials.
@@
We apply the PPO algorithm (Schulman et al., 2017) with an Actor & Critic networks and a Clip objective function to efficiently learn our MACB.
  • Actor & Critic Definition. To model the selection of reflections and evaluate the value of the states, we implement a policy network parameterized by MLPs to define our selector’s Actor $\pi_{\bm{{\theta}}}$ and Critic $V_{\bm{\psi}}$, which maps the environmental space $\mathcal{Z}$ to the action space $\mathcal{A}$ and reward function $R$ respectively:
    (7) $\mathrm{Actor\ }\pi_{\bm{\theta}}(\bm{z}):\bm{a}$ $=\mathrm{softmax}(\bm{\theta}\cdot\bm{z})$
    $\mathrm{Critic\ }V_{\bm{\psi}}(\bm{z}):r$ $=\bm{\psi}\cdot\bm{z}$
    where $\bm{\theta}$ and $\bm{\psi}$ are the learnable weights of Actor and Critic respectively.
  • Clip Objective. During the training process, the selector’s Actor uses the $\epsilon$-greedy ($\epsilon=0.1$) strategy to explore the environment, with probability $\epsilon$ to take a random action while with probability ($1-\epsilon$) to exploit the learned policy $\pi_{\bm{\theta}}$. We adopt the clip objective function for PPO training to maximize the reward selection as follows:
@@
During inference, we select the reflection memories with the highest confidence from the Actor, exploit the reflection $Ref_{i}$ with the highest $Imp$ score in the memory, and make a recommendation according to Eq. 2 by replacing the prompt $P$ with the reflection $Ref_{i}$.

### v2

@@ 3.3. Offline: Multi-Perspective Reflections
As illustrated on the left of Fig. 3, MoRE aims at exploiting users’ historical interaction to generate reflections from the perspectives of explicit preference, implicit preference, and collaborative signals. We achieve this goal through the following reflectors:
Explicit Preference (EP) Reflector captures user $u$’s explicit preferences by leveraging the LLM to reflect on the user’s interaction history $\mathcal{S}_{u}$, the candidate items $\mathcal{C}_{u}^{t}$, and the discrepancy between the predicted ranking $\hat{\mathcal{O}}_{u}^{t}$ generated from LLM-based recommender and the ground-truth $v_{u}^{t}$.
The reflection $Ref_{\mathrm{EP},u}$ from the perspective of explicit preference can be formulated as follows:
(6) $Ref_{EP,u}=\mathrm{LLM}_{\mathrm{REF}}(P_{EP}(u,\mathcal{S}_{u}^{t-1},\mathcal{C}_{u}^{t},\hat{\mathcal{O}}_{u}^{t},v_{u}^{t})),$
where $P_{{EP}}$ denotes the prompt specifically designed to guide the LLM in reflecting explicit user preferences, defined as follows:
EP reflection prompt $P_{EP}$ You are a reflector for an LLM-based recommender system, understanding the explicit preferences embodied in the user’s historical sequence and analyzing the areas for improvement in past recommendation attempts to provide reflections for the future. Explicit preferences are derived from an analysis of recent tendencies reflected in the sequence of item titles and descriptions within the user’s history. You need to:
1. Analyze the history with associated text to identify explicit preferences. 2. Analyze the logic and rationale behind past recommendation attempts. 3. Examine potential shortcomings in the past and provide suggestions for improvement.
User’s historical sequence with related description: $\langle REP_{EP}(S_{u}^{t-1})\rangle$. Candidates: $\langle REP_{EP}(C_{u}^{t})\rangle$. Past recommendation attempts: $\langle REP_{EP}(\hat{O}_{u}^{t})\rangle$. User new interaction (if any): $\langle REP_{EP}(o_{u}^{t})\rangle$. Historical reflection demonstrations (if any): $\langle DEMO\rangle$. Your reflection:
Please note that in EP reflector, item titles with descriptions are used to represent $\mathcal{S}_{u}^{t-1},\mathcal{C}_{u}^{t},\hat{\mathcal{O}}_{u}^{t},$ and $v_{u}^{t}$ through $REP_{EP}()$.
Implicit Preference (IP) Reflector is designed to capture the preference embedded in the other item attributes (e.g., item brands). Similar to the EP reflector, the IP reflector takes the interaction history $\mathcal{S}_{u}$ of user $u$, the candidate item list $\mathcal{C}_{u}$, the predicted ranking list $\hat{\mathcal{O}}_{u}$, and the ground-truth item $v_{u}^{t}$ as input, and produce the reflection through a frozen LLM:
(7) $Ref_{IP,u}=\mathrm{LLM}_{\mathrm{REF}}(P_{IP}(u,\mathcal{S}_{u}^{t-1},\mathcal{C}_{u}^{t},\hat{\mathcal{O}}_{u}^{t},v_{u}^{t})),$
where $P_{{IP}}$ denotes the prompt for guiding the LLM in reflecting implicit user preferences, defined as follows:
IP reflection prompt $P_{IP}$ You are a reflector for an LLM-based recommender system, understand the implicit preferences embodied in the user’s historical sequence and analyze the areas for improvement in past recommendation attempts to provide reflections for future. Implicit preferences are reflected through the associations shown in subsequences of item attributes within the user’s history, such as subsequences of brands, styles, functions, features, etc. Requirements:
1. Analyze the logic and rationale behind past recommendation attempts.
2. Focus on attribute subsequences within history to analyze potential associations and causality.
3. Examine possible shortcomings in the past and provide suggestions.
Historical attribute seq: $\langle REP_{IP}(S_{u}^{t-1})\rangle$. Candidates: $\langle REP_{IP}(C_{u}^{t})\rangle$. Past recommendation attempts: $\langle REP_{IP}(\hat{O}_{u}^{t})\rangle$. User new interaction (if any): $\langle REP_{IP}(o_{u}^{t})\rangle$. Historical reflection demonstrations (if any): $\langle DEMO\rangle$. Your reflections:
In IP reflector, the attributes of items are used to represent $\mathcal{S}_{u}^{t-1},\mathcal{C}_{u}^{t},\hat{\mathcal{O}}_{u}^{t},$ and $v_{u}^{t}$ through $REP_{IP}()$.
Collaborative Filtering (CF) Reflector aims to integrate patterns and similarities across user interaction sequences into LLM-based recommender systems. Unlike approaches that introduce new token embeddings into the vocabulary (Zheng et al., 2024; Liao et al., 2024) or insert special sequences into prompts (Zhang et al., 2024a), we leverage a pre-trained collaborative filtering model $\mathcal{M}_{u}$ to incorporate CF ratings into the LLM reflection process. Specifically, the CF reflector takes as input the target user $u$’s interaction history $\mathcal{S}_{u}^{t-1}$, the candidate item set $\mathcal{C}_{u}^{t}$, the predicted ranking $\hat{\mathcal{O}}_{u}^{t}$, and the ground-truth item $v_{u}^{t}$, and generates reflections using a frozen LLM:
(8) $Ref_{CF,u}=\mathrm{LLM}_{\mathrm{REF}}(P_{CF}(u,\mathcal{S}_{u}^{t-1},\mathcal{C}_{u}^{t},\hat{\mathcal{O}}_{u}^{t},v_{u}^{t})),$
where $P_{{CF}}$ is the prompt designed to guide the LLM in reflecting CF signals. The prompt structure is detailed as follows:
CF reflection prompt $P_{CF}$ You are a reflector for an LLM-based recommender system, utilizing a Collaborative Filtering (CF) model to obtain items’ CF signals and analyzing potential improvement in past recommendation attempts for future suggestions. The sequence of historical items’ CF ratings reflects patterns and similarities across user sequences. Your step-by-step task:
1. Analyze historical CF ratings to identify preferences, trends, or interest shifts.
2. Evaluate past recommendation effectiveness by comparing suggested items with actual user interactions.
3. Analyze potential shortcomings in past recommendation attempts and provide suggestions for improvement.
Items presented with CF ratings: $\langle REP_{\mathcal{M}_{u}}(S_{u}^{t-1},C_{u}^{t},\hat{O}_{u}^{t},o_{u}^{t})\rangle$ Effective historical reflection demonstrations (if any): $\langle DEMO\rangle$. Your reflection:
In $P_{{CF}}$, $\mathcal{S}_{u}^{t-1},\mathcal{C}_{u}^{t},\hat{\mathcal{O}}_{u}^{t},$ and $v_{u}^{t}$ are represented by their corresponding CF ratings generated by $\mathcal{M}_{u}$ which is denoted as ¡$REP_{\mathcal{M}_{u}}$¿.
@@
For each user, we maintain three distinct memory banks, each dedicated to storing the generated reflections from one of the three perspectives (Explicit Preference, Implicit Preference, and Collaborative Filtering respectively). Each memory bank for a user stores a collection of past reflections for that user from the specific perspective. These memory banks are guided by the meta-reflector to improve the reflections, as will be elaborated in the following section.
3.4. Offline: Reflection Memory Maintenance
@@
Reflection Improvement Effect is defined as the difference in LLM recommendation performance on the validation set, with or without the reflections. In other words, if incorporating reflections improves the LLM’s performance, it indicates that the reflection is effective for recommendation. The greater the performance improvement, the more effective the reflection. If the reflection improvement effect of a reflection exceeds a preset threshold $h$, we regard it as an effective one.
Formally, we formulate the reflection improvement effect as:
(9) $Imp$ $=\mathrm{Metric}(\hat{\mathcal{O}}_{u}^{\mathrm{REF}})-\mathrm{Metric}(\hat{\mathcal{O}}_{u}^{\mathrm{REC}}),$
$\hat{\mathcal{O}}_{u}^{\mathrm{REF}}$ $=\mathrm{LLM}_{\mathrm{REC}}(P_{\mathrm{REC}}(u,\mathcal{S}_{u}^{t-1},\mathcal{C}_{u}^{t},Ref_{i,u})),$
$\hat{\mathcal{O}}_{u}^{\mathrm{REC}}$ $=\mathrm{LLM}_{\mathrm{REC}}(P_{\mathrm{REC}}(u,\mathcal{S}_{u}^{t-1},\mathcal{C}_{u}^{t})),$
where $Ref_{i,u}\in\{Ref_{\mathrm{EP},u},Ref_{\mathrm{IP},u},Ref_{\mathrm{CF},u}\}$, $\mathrm{Metric}$ denotes the recommendation metric, e.g., NDCG@10.
@@
(2) Group Level: We use a CF model to construct embeddings for all users and then cluster them using the K-means++ algorithm (Bahmani et al., 2012). We then choose the reflections that provide the greatest average improvement for users within the same cluster, ensuring that users in the same cluster share the same reflections.
@@
All of the three strategies are greedy, which may lead to local optima. To mitigate this issue, we propose using the improvement effect brought about by each reflection as the sampling probability. Reflections sampled from the filtered set serve as in-context learning demonstrations. These demonstrations are then used to task the three reflectors with generating improved reflections in subsequent iteration steps, enabling the LLM to self-improve its reflection quality.
@@
3.5. Online: Reflection Perspective Selection
@@
The second objective of the meta-reflector is to model preference shifts in SeqRec. Given three memory banks for each user to collect effective reflections from three different perspectives, we require the meta-reflector to select one of these reflections for recommendation to achieve this goal. We formulate the decision-making process as a Multi-Armed Contextual Bandit (MACB) problem (Li et al., 2010), and adopt the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017) to efficiently learn to select the reflections. Formally, we define MACB as a tuple $(\mathcal{Z},\mathcal{A},\mathcal{R})$, where $\mathcal{Z},\mathcal{A},\mathcal{R}$ denote the state space, action (arm) space, and reward function, respectively.
@@
(1) state $\bm{z}\in\mathcal{Z}$ represents recommendation context, constructed by concatenating CF-derived embeddings: $\bm{z}=[emb_{u},emb_{v}]$ where $emb_{u}$ is the user embedding and $emb_{v}$ the average interaction embedding from $\mathcal{S}_{u}^{t-1}$.
@@
(2) Action $\bm{a}\in\mathcal{A}$ is defined as the selection of a reflection from three memory banks. We represent $\bm{a}$ as a 3-dimensional vector, with each dimension corresponding to an arm, each arm representing a reflection memory.
@@
(3) Reward is defined to assess the recommendation improvement brought by the reflections. We reuse the measure in Eq.(9) and represent the reward as $R(\bm{z,a})=Imp$. Since the reflection has been assessed in Sec. 3.3 we do not need to call the LLM here. Additionally, due to the sampling discussed in Sec. 3.4, $R(\bm{z,a})$ can be both positive and negative, ensuring training efficiency.
@@
(4) Replay Buffer is designed to facilitate efficiency in policy optimization. We represent it as $D=(\bm{z},\bm{a},R(\bm{z},\bm{a}),\bm{z}^{\prime})$ to store the tuples of observed state, action, reward, and next state. With the records in the replay buffer, we can refine successful policies and learn from erroneous trials.
Table 1. Statistics of the 3 pre-processed datasets.
Dataset #Users #Items Avg. len Sparsity
Arts 55970 22612 8.80 99.96%
Games 55145 17287 9.01 99.95%
Instruments 27404 10450 8.41 99.92%
Table 2. Recommendation performance comparison. All LLM-based methods utilize Llama-3 and ensure an identical candidate set for a fair comparison. The best and the second-best performances are denoted in bold and underlined fonts, respectively. “N@K” is short for “NDCG@K”. The “Imp.” indicates the percentage improvement of MoRE over the best performances from baselines. “-” means that the baseline can only output a top-1 prediction and cannot rank or recall multiple items. denotes MoRE performs significantly better than baselines based on two-tailed paired t-test with Bonferroni correction ($p<0.05$).
Methods Arts Games Instruments
HR@1 HR@5 HR@10 N@5 N@10 HR@1 HR@5 HR@10 N@5 N@10 HR@1 HR@5 HR@10 N@5 N@10
Caser 0.103 0.2555 0.3603 0.1823 0.2161 0.0899 0.3024 0.4397 0.1997 0.2440 0.1349 0.2931 0.3968 0.2193 0.2527
FDSA 0.1211 0.2536 0.3670 0.1875 0.2241 0.1196 0.3073 0.4466 0.2150 0.2597 0.1403 0.2842 0.3905 0.2157 0.2497
BERT4Rec 0.0896 0.2402 0.3613 0.1680 0.2072 0.0899 0.2796 0.4160 0.1861 0.2299 0.1233 0.2618 0.3789 0.1954 0.2326
GRU4Rec 0.1001 0.2431 0.3708 0.1722 0.2129 0.0929 0.3192 0.4644 0.2083 0.2554 0.1340 0.3021 0.4021 0.2201 0.2520
SASRec 0.1850 0.3570 0.4280 0.2788 0.3016 0.1900 0.2915 0.4348 0.1898 0.2369 0.1832 0.3327 0.3701 0.2594 0.2809
LLM4RS 0.0543 0.1087 0.1277 0.0826 0.0887 0.0662 0.1976 0.2233 0.1338 0.1421 0.0330 0.0900 0.1188 0.0624 0.0717
LLMRank(CoT) 0.1725 0.3337 0.4061 0.2595 0.2828 0.1611 0.3508 0.4496 0.2607 0.2925 0.1457 0.2904 0.3709 0.2202 0.2461
LC-Rec 0.2145 0.3768 0.4309 0.3006 0.3181 0.2599 0.4140 0.4646 0.3426 0.3593 0.1592 0.3296 0.3920 0.2486 0.2691
Re2LLM 0.2736 0.4252 0.4938 0.3535 0.3757 0.2223 0.4180 0.5198 0.3231 0.3559 0.2011 0.3521 0.4424 0.2787 0.3074
MoRE 0.2898 0.4376 0.5043 0.3705 0.3922 0.2569 0.4773 0.5731 0.3647 0.3957 0.2163 0.3824 0.4674 0.2981 0.3251
Imp. 5.92% 2.92% 2.13% 4.81% 4.39% -1.15% 14.19% 10.25% 6.45% 10.13% 7.56% 8.61% 5.65% 6.96% 5.76%
@@
We apply the PPO algorithm (Schulman et al., 2017) with an Actor & Critic network and a Clip objective function to efficiently learn MACB.
(1) Actor & Critic Definition. To model the selection of reflections and evaluate the value of the states, we implement a policy network parameterized by MLPs to define our selector’s Actor $\pi_{\bm{{\theta}}}$ and Critic $V_{\bm{\psi}}$, which maps the environmental space $\mathcal{Z}$ to the action space $\mathcal{A}$ and reward function $R$ respectively:
(10) $\mathrm{Actor\ }\pi_{\bm{\theta}}(\bm{z}):\bm{a}$ $=\mathrm{softmax}(\bm{\theta}\cdot\bm{z}),$
$\mathrm{Critic\ }V_{\bm{\psi}}(\bm{z}):r$ $=\bm{\psi}\cdot\bm{z},$
where $\bm{\theta}$ and $\bm{\psi}$ are the parameters of Actor and Critic respectively.
(2) Clip Objective. During the training process, the selector’s Actor uses the $\epsilon$-greedy ($\epsilon=0.1$) strategy to explore the environment, with probability $\epsilon$ to take a random action while with probability ($1-\epsilon$) to exploit the learned policy $\pi_{\bm{\theta}}$. We adopt the clip objective function for PPO training to maximize the reward selection as follows:
@@
During inference, we select the reflection memories with the highest confidence from the Actor, exploit the reflection $Ref_{i,u}$ with the highest $Imp$ score in the memory, and make a recommendation according to Eq.(2) using the prompt $P_{\mathrm{REC}}$.
4. Experiment
In this section, we conduct extensive experiments to answer the following research questions: RQ1: How does MoRE compare to various existing baselines in terms of recommendation performance? RQ2: How effective are the multi-perspective reflections and reflection perspective selection in MoRE? RQ3: For reflection memory maintenance, how effective are the refining and iteration strategies? RQ4: Does MoRE have an advantage in terms of training cost?
Table 3. Comparison of HR@1 Performance with LLM baselines that inherently only provide top-1 prediction.